### v2

@@ Abstract
keywords:
Human walking locomotion, Granular terrain, Gait adaptation
[…]
organization=Department of Mechanical and Aerospace Engineering
Rutgers University,city=Piscataway, state=NJ 08854, country=USA
@@ 1 Introduction
The biomechanical study of human locomotion has been a fundamental aspect of understanding human physiology and developing assistive devices and wearable robotic systems. However, the majority of gait analysis research and open-source datasets were performed on level, solid ground (Camargo et al., 2021). This focus has yielded significant insights into the basic biomechanics of gait but has not encapsulated the full spectrum of challenges and variables present during everyday ambulation. When it comes to real-world locomotion, navigating through granular terrains such as sand or small pebbles can be particularly challenging due to their constantly yielding and shifting nature (Kowalsky et al., 2021).
@@
Several research works have provided a comprehensive understanding of how different terrain conditions affect human walking locomotion. The metabolic cost of walking on compliant substrates was investigated in Grant et al. (2022), and the study provided a foundation for understanding the energy demands of walking on yielding terrains such as sand. The work of Zamparo et al. (1992) specifically examined the energy cost of walking and running on sand and highlighted the increased physical effort required on yielding surfaces. These studies underline the importance of terrain in influencing gait mechanics and energy efficiency. The research work on human gait in natural settings (Kowalsky et al., 2021) and on diverse surfaces such as forest terrains (Holowka et al., 2022) provided a comprehensive picture of the interaction between terrain type, gait parameters, and energy expenditure. Specific adaptations to walking on uneven terrains such as unanticipated steps were observed and corresponding lower limb bio-mechanical response was analyzed in Panizzolo et al. (2017). Another comprehensive overview of the terrain impacts on human sprint locomotion demonstrated the effects of different surfaces such as natural and artificial turf, and sand. The study in Xu et al. (2015) discussed the influence of deformation height on estimating the center of pressure (COP) during walking on level ground and sand.
@@
Recent advancements in assistive technologies have significantly contributed to the understanding of human-terrain interaction during walking locomotion. Developments in exoskeleton control for uneven terrains in  Jatsun et al. (2018) and the integration of human-in-the-loop control in soft exosuits on different terrains in Li et al. (2022) have enhanced activity recognition and response adaptability. The use of inertial measurement unit (IMU)-based algorithms for foot-ground contact detection (Kim and Lee, 2017) and terrain topography detection (Knuth and Groves, 2023) illustrates the complexity of walking dynamics on varied surfaces. These technologies play a crucial role in real-time walking gait phase estimation, especially on uneven terrain (Medrano et al., 2023). Additionally, studies focusing on the asymmetry of ground contact during running on sand (Dewolf et al., 2019) and the analysis of ground reaction forces and muscle activity in walking across different surfaces (Jafarnezhadgero et al., 2019) have extended the understanding of the biomechanical adaptations required for locomotion on diverse terrains. Moreover, Grant (2023) provided a holistic view of how human gait and energetics are altered when traversing substrates of varying compliance.
@@
All the above-mentioned research emphasizes the complex interplay between human biomechanics and varying terrain types. Comprehensive joint-level biomechanics analysis on locomotion over such yielding terrain however remains scant. Recognizing this deficit, this study introduces a detailed analysis of human locomotion on two distinct terrains: solid ground and sand. The methodology involves collecting and analyzing a novel dataset of biomechanical data from $20$ able-bodied adults. A motion capture system coupled with ground reaction force (GRF) measurements is used to capture both kinematic and kinetic aspects of walking locomotion. We comprehensively analyze joint-level biomechanics, including GRFs and joint torque profiles, in environments that replicate natural walking conditions on both solid and yielding sand surfaces.
@@
The main contributions of this work are twofold. First, an in-depth analysis of the biomechanical characteristics of human walking on granular terrains is presented. To the authors’ best knowledge, there is no reported result on GRFs and joint torque profiles and corresponding comparisons for humans walking both on sand and solid ground. This study fills such a knowledge gap and offers new insights into the biomechanical strategies employed by individuals when navigating through granular terrains—a scenario frequently encountered in daily life but seldom studied in detail. The biomechanical analysis in this work not only contributes to enhancing the understanding of human gait mechanics but also serves as a foundation for the development of assistive devices. Second, this work provides an experimental dataset that integrates synchronized motion capture and GRF data. This dataset is expected to be a valuable resource for further biomechanics and wearable sensor data research across varied environmental settings.
@@ 2.1 Experimental Setup and Protocol
In this study, twenty able-bodied healthy subjects ($14$ males and $6$ females, age: $24.8\pm 4.0$, height: $171.7\pm 9.5$ cm, weight: $74.5\pm 16.1$ kg) participated in experiments and are labeled as S1 to S20. All participants are self-reported to be in a good health condition. An informed consent form was signed by each subject and the experimental protocol was approved by the Institutional Review Board (IRB) at Rutgers University.
Figure 1 shows the experimental setup in a laboratory. A three-segment walkway ($7.5$ m long, $0.76$ m wide, and $0.14$ m high) was constructed as shown in the figure. Segment $1$ was built with reinforced plywood with one embedded force plate (model ACG-O from Advanced Mechanical Technology, Inc., Watertown, MA) to replicate a solid ground condition. Segment $2$ was a sand box filled with sand as a common granular terrain condition. Another force plate (from Bertec Corporation.) was buried $14$ cm beneath the sand surface to capture the GRFs. The GRF measurements collected within Segment $2$ were calibrated by a correlation for the sand layer thickness. The correlation curves can be found in A. Segment $3$, made of the same material as segment one, was connected to segment two to ensure subjects could comfortably leave segment two without changing their gaits.
Figure 1: Experimental platform design and setup inside a laboratory.
Starting at Segment $1$ of the walkway, each subject was instructed to walk in the preferred manner. For the first $5$ trials, the subject was required to walk back and forth on the walkway to get familiar with walking with the sensing devices and the terrain condition of the walkway. After that, another $6$-$8$ trials of one-directional walking experiments were conducted. The sand surface was flattened before each trial.
2.2 Sensor Measurements and Data Collections
For the walking motion, a motion capture system ($10$ Vantage cameras, Vicon Motion Systems Ltd.) was used to collect lower limb motion information with the sampling frequency $100$ Hz. A total of $18$ reflective markers were carefully attached firmly to each subject. Figure 2 shows the specific schematic of Vicon marker positions and their name labels. “L” and “R” represent “left” and “right”, respectively. The subjects were required to wear a pair of swimming boots for the free movement of foot navicular joints. Meanwhile, the GRF measurements were synchronized via hardware and software setup with the motion data at $1000$ Hz.
@@ 2.3 Data Processing
Major gait events such as heel strike and toe-off were defined through motion data to identify the stance phases (i.e., single and double stance phase, respectively) of the gait cycle. The walking gait progression, namely, the joint angle, was normalized by a defined phase variable (between 00 and $1$) according to the strike events. For clear presentation, the GRFs were normalized the weight of the subject and presented as the functions of the stance variable which was normalized between the heel-strike and toe-off.
The measurements and calculations of interest are joint angles of hip, knee, and ankle in the sagittal plane of the human walking locomotion, namely, the $ZOX$ plane shown in Fig. 1. Likewise, we only presented the longitudinal and vertical forces in the $X$-and $Z$-direction, respectively. Joint moments in the sagittal plane were also computed through an inverse dynamics model given the GRFs applied on the foot; see the detailed description in B. Other kinematic variables such as strike length/width, stance/swing time, vertical variation of center of mass (COM), and average walking velocity highlighted significant variations in the gait patterns between solid ground and sand, and they were included and reported in the study. The data processing was done by using custom algorithms in MATLAB software (Mathworks Inc., Natick, MA). A one-way repeated measures analysis of variance (ANOVA) was conducted to evaluate the impact of the terrains on the above-mentioned metrics.
@@ 3.1 Kinematic Analysis Results
Strike Differences: Figure 3 illustrates stride profiles among all the subjects. Stride length was found to be significantly longer (i.e., $20$%) on sand, with an accompanying increase of $17$% in stride width. This gait adaptation is likely due to the deformable nature of the sandy terrain. The COM vertical variation during each step on the sand was also observed to increase by $6$%, which was possibly due to the extra effort from the joints when subjects walked on the sand. Walking on sand also showed an obvious increase regarding the swing time and stance time. Overall, subjects tended to walk faster on sand than on solid ground. The ratio between average walking speed on solid ground versus that on sand was about $0.94$.
Figure 3: The distribution of step parameters for all participants combined (n = $20$) while walking on the two different terrains: solid ground (blue) and sand (yellow). The comparison was among the stride length, stride width, COM variation, normalized walking velocity (all normalized with respect to subjects’ heights), swing, and stance time. Data included all strides for individual trials. The label *** indicates a significant ($p<0.05$) difference among trials.
Joint Angles: Differences in joint angles, particularly in the lower extremities, were observed, and this indicates a distinct adjustment in walking strategy on sand compared to that on solid ground. Figure 4 shows the detailed profiles for the hip, knee, and ankle joints throughout the gait cycle. There was no clear difference regarding the joint angle pattern. Only slightly larger magnitudes were observed in the hip flexion/extension and knee flex/extension for walking on sand than those on solid ground. However, for ankle dorsi/plantarflexion, this difference did not clearly occur in the gait pattern.
@@ 3.2 Kinetic Analysis Results
The kinetic analysis provides insights into how walking on different terrains affects the forces exerted by and on the body:
GRFs: Figures 5 and 5 show the longitudinal and vertical GRFs normalized by the subject’s weight, respectively. The shaded region indicated the one-standard variation. It is observed that the magnitude of the maximum longitudinal force $F_{x}$ on sand is relatively smaller than that on solid ground in both forward and backward directions. Furthermore, this noticeable disparity appears in the toe pushing off stage. On solid ground, the average pushing force is approximately four times larger than that on sand. However, regarding the vertical force $F_{z}$, sandy terrain might provide more supporting force during the heel strike and foot rolling stages, compared with the solid ground. Moreover, the variation of the normal force on the sand is also more significant than that on the solid ground. It is also noticeable that there is a slight difference with the double-hump pattern of the vertical force $F_{z}$ throughout the contact phase. The magnitudes of two humps on the solid ground appear almost equivalent. Nevertheless, for the sandy terrain, the first hump of the force is slightly higher than the second one in the late contact phase. Similar results were also reported on the level sand in Xu et al. (2015).
Joint Moments: Figure 6 illustrates the joint moments (Nm/kg) in the sagittal plane normalized by the subject’s mass. The inverse dynamics calculation model used in this study was discussed in Shamaei et al. (2013) and the detailed description is presented in B. From the comparison results, we observe that for able-bodied locomotion, the joint moments kept consistent bilaterally both on sandy and solid terrain.
@@ 4 Discussion
Strike and COM: Different from the experiment setup designed in Svenningsen et al. (2019), the subjects in this study were asked to walk across two different terrains sequentially from the solid ground to the sand box, allowing each subject to keep the preferred comfortable walking pace. Our findings suggest that walking strategies adapt significantly when transitioning from solid ground to sand. Participants tended to walk faster on sand than on solid ground; see the velocity ratio result in Fig. 3. The extended stride length and width observed on the sand were likely compensatory mechanisms to maintain balance on the yielding surface. Additionally, the increased COM vertical variation indicates an adaptive response to optimize energy efficiency and maintain stability on the yielding terrain, which was also found for the mid-stance phase ($21\%-60\%$ gait cycle) in Svenningsen et al. (2019). These adaptations are pivotal for navigating through granular substrates, where it is challenging for gait strategies to navigate under solid ground.
@@
Kinematics: The joint angles, especially in the lower extremities, demonstrate distinct variations between solid and sand terrains. Walking on the sandy terrain results in greater dorsiflexion of the ankle in the early stance phase but a similar peak amplitude of the ankle plantar flexion near toe-off. This is because the heel would experience an intrusion into the sand and then the forefoot was prevented from plantarflexion at the beginning of the stance phase; see the small portion of ankle plantarflexion on the solid ground shown as the blue dashed line in Fig. 4. Furthermore, a relatively larger variability of the ankle dorsi/plantarflexion on the sand was found compared with the hip flexion/extension and knee flex/extension. This represents the nature of complexity and instability of deformable terrain locomotion. Similarly, greater flexion is observed in the hip and knee profile. The high flexion in the toe-off and swing phase can be a compensation for the loss of momentum during the stance phase for the deformable sand.

### v3

@@ 1 Introduction
The biomechanical study of human locomotion has been a fundamental aspect of understanding human physiology and developing assistive devices and wearable robotic systems. However, the majority of gait analysis research and open-source datasets were performed on level, solid ground [1]. This focus has yielded significant insights into the basic gait biomechanics but has not encapsulated the full spectrum of challenges and variables present during everyday ambulation. When it comes to real-world locomotion, navigating through granular terrains such as sand can be particularly challenging due to their constantly yielding and shifting nature [2]. Sand and other yielding surfaces are prevalent in everyday life, including beaches, playgrounds, and off-road trails [3]. The distinct properties of sand, such as its constantly shifting and deformable nature, present unique challenges for maintaining stability and efficiency during locomotion [4].
@@
Several research works have provided a comprehensive understanding of how different terrain conditions affect human walking locomotion. The metabolic cost of walking on compliant substrates was investigated in [5], and the study provided a foundation for understanding the energy demands of walking on sand. The work of [4] specifically examined the energy cost of walking and running on sand and highlighted the increased physical effort required on yielding surfaces. The mechanical work and energetic cost during walking and running on sand and on a hard surface were compared in [6]. In [7], the authors quantitatively compared the speed, temporal segmentation, and variability of walking on solid ground and sand. These studies underline the importance of terrain in influencing gait mechanics and energy efficiency.
The research work on human gait in natural settings [2] and on diverse surfaces [8] provided a comprehensive picture of the interaction between terrain type, gait parameters, and energy expenditure. Specific adaptations to walking on uneven terrains such as unanticipated steps were observed and corresponding lower limb biomechanical response was analyzed in [9]. Another comprehensive overview of the terrain impacts on human sprint locomotion demonstrated the effects of different surfaces such as natural and artificial turf, and sand. The study in [10] discussed the influence of deformation height on estimating the center of pressure (COP) during walking on level ground and sand, showing that the magnitude of COP changes in the anterior-posterior (AP) and medial-lateral (ML) directions differed in both level and cross-slope conditions.
Figure 1: Experimental platform design and setup inside a laboratory for human walking on solid and sand terrains.
@@
Recent advancements in assistive technologies have significantly contributed to the understanding of human-terrain interaction during walking locomotion. Developments in exoskeleton control for uneven terrains in [11] and the integration of human-in-the-loop control in soft exosuits on different terrains in [12] have enhanced activity recognition and response adaptability. The use of inertial measurement unit (IMU)-based algorithms for foot-ground contact detection [13, 14] and terrain topography detection [15] illustrates the complexity of walking dynamics on varied surfaces. These technique developments play a crucial role in real-time walking gait phase estimation, especially on uneven terrain [16]. Additionally, studies focusing on the asymmetry of ground contact during running on sand [17] and the analysis of ground reaction forces and muscle activity in walking across different surfaces [18, 19] have extended the understanding of the biomechanical adaptations required for locomotion on diverse terrains. Moreover, [20] provided a holistic view of how human gait and energetics are altered when traversing substrates of varying compliance.
@@
Understanding the biomechanics of walking on sand has potential implications for various applications, such as rehabilitation, biped/quadruped robot control, and human exoskeleton assistance. The knowledge gained from studying human locomotion on sand can be applied to improve the control and stability of bipedal robots, as well as to enhance the performance of exoskeletons in assisting human movement on diverse surfaces [21]. Therefore, evaluating the biomechanics of walking on sand is essential for advancing our understanding of human locomotion in diverse terrains and developing assistive devices and robotic systems that can adapt to these common environments. All the above-mentioned research emphasizes the complex interplay between human biomechanics and varying terrain types. Comprehensive joint-level biomechanics analysis on locomotion over such yielding terrain however remains scant. Recognizing this deficit, this study introduces a detailed analysis of human locomotion on two distinct terrains: solid ground and sand. The methodology involves collecting and analyzing a novel dataset of biomechanical data from $20$ able-bodied adults. A motion capture system coupled with ground reaction force (GRF) measurements is used to capture both kinematic and kinetic aspects of walking locomotion. We comprehensively analyze joint-level biomechanics, including GRFs and joint torque profiles, in environments that replicate natural walking conditions on both solid and yielding sand surfaces.
@@
The main contributions of this work are twofold. First, an in-depth analysis of the biomechanical characteristics of human walking on granular terrains is presented. To the authors’ best knowledge, there is no reported result on GRFs and joint torque profiles and corresponding comparisons for humans walking both on sand and solid ground. This study fills such a knowledge gap and offers new insights into the biomechanical strategies employed by individuals when navigating through granular terrains. The biomechanical analysis in this work not only contributes to enhancing the understanding of human gait mechanics but also serves as a foundation for the development of assistive devices. Second, this work provides an experimental dataset that integrates synchronized motion capture and GRF data. This dataset is expected to be a valuable resource for further biomechanics and wearable sensor data research across varied environmental settings.
@@ 2.1 Experimental Setup and Protocol
In this study, twenty able-bodied healthy participants ($14$ males and $6$ females, age: $24.8\pm 4.0$ years old, height: $171.7\pm 9.5$ cm, weight: $74.5\pm 16.1$ kg) participated in experiments and are labeled as S1 to S20. All participants are self-reported to be in a good health condition. An informed consent form was signed by each participant and the experimental protocol was approved by the Institutional Review Board (IRB) at Rutgers University.
Figure 1 shows the experimental setup in a laboratory. A three-segment walkway ($7.5$ m long, $0.76$ m wide, and $0.14$ m high) was constructed as shown in the figure. Segment $1$ was built with reinforced plywood with one embedded force plate (model ACG-O from Advanced Mechanical Technology, Inc., Watertown, MA) to replicate a solid ground condition. Segment $2$ was a sandbox filled with sand to represent a common granular terrain condition. A force plate (from Bertec Corporation) was buried $14$ cm beneath the sand surface to capture the ground reaction forces (GRFs). Although the work in [10] studied the influence of deformation height on the COP, the dissipation of GRFs was not considered. To account for this, the GRF measurements collected within Segment $2$ were calibrated using a correlation based on the sand layer thickness. The calibration process and resulting correlation curves can be found in Appendix A. Segment $3$, made of the same material as segment one, was connected to segment two to ensure participants could comfortably leave segment two without changing their gaits.
The experimental setup is designed to study gait adaptations and changes when traversing different terrains. While this setup differs from previous studies that typically analyze gait on a single, uniform surface for several consecutive strides, this design captures the immediate adaptations and biomechanical adjustments that occur during real-world terrain transitions. To ensure meaningful results, we ignored the first few strides on each terrain and focused our analysis on a single, representative step from each trial, allowing participants to adjust to the new surface conditions before recording the relevant data.
Starting at Segment $1$ of the walkway, each participant was instructed to walk in the preferred manner. For the first $5$ trials, the participant was required to walk back and forth on the walkway to get familiar with walking with the sensing devices and the terrain condition of the walkway. After that, another $6$-$8$ trials of one-directional walking experiments were conducted. The sand surface was flattened before each trial.
For the walking motion, a motion capture system ($10$ Vantage cameras, Vicon Motion Systems Ltd.) was used to collect lower limb motion information with the sampling frequency $100$ Hz. A total of 18 reflective markers were firmly attached to each participant using straps. Figure 2 shows the specific schematic of Vicon marker positions and their corresponding labels. “L” and “R” denote “left” and “right”, respectively. Using the marker data, a customized lower-limb model was constructed within the Vicon Nexus software for subsequent analysis. The participants were required to wear a pair of swimming boots for the free movement of foot navicular joints. Meanwhile, the GRF measurements were synchronized via hardware and software setup with the motion data at $1000$ Hz.
@@
Figure 3: The distribution of step parameters for all participants combined (n = $20$) while walking on the two different terrains: solid ground (blue) and sand (yellow). The comparison was among the stride length, stride width, COM variation, walking velocity (all normalized with respect to participants’ heights), swing, and stance time. Data included all strides for individual trials. The label * indicates a significant ($p<0.05$) difference among trials.
@@ 2.2 Data Processing
Major gait events such as heel strike and toe-off were defined through motion data to identify the stance phases (i.e., single and double stance phases, respectively) of the gait cycle. The GRFs in the horizontal and vertical directions were smoothed by a moving average filter. The walking gait progression, namely, the joint angle, was normalized by a defined phase variable (between 00 and $1$) according to the strike events. For clear presentation, the GRFs were normalized the weight of the participant and presented as the functions of the stance variable which was normalized between the heel-strike and toe-off.
The measurements and calculations of interest are the angles of the hip, knee, and ankle joints in the sagittal plane of the human walking locomotion, namely, the $ZOX$ plane shown in Fig. 1. Likewise, we only presented the longitudinal and vertical forces in the $X$-and $Z$-direction, respectively. Joint moments in the sagittal plane were also computed through an inverse dynamics model given the GRFs applied on the foot; see the detailed description of the joint moments calculation in Appendix B. Other kinematic variables such as stride length/width, stance/swing time, vertical variation of center of mass (COM), and average walking velocity highlighted significant variations in the gait patterns between solid ground and sand, and they were included and reported in the study. The data processing was done by using custom algorithms in MATLAB software (Mathworks Inc., Natick, MA). A paired sample t-test was conducted to evaluate the impact of the terrains on the above-mentioned metrics.
@@ 3.1 Kinematic Analysis Results
Figure 3 illustrates stride profiles among all the participants. Compared with that on solid ground, stride length was found to be significantly longer (i.e., $20$%) on sand, with an accompanying increase of $17$% in stride width. Statistical analysis revealed that walking on sand resulted in a significantly larger normalized stride length ($0.74\pm 0.08$) compared with on solid ground ($0.663\pm 0.19$) ($p<0.05$; $d=0.53$). Similar changes in normalized stride widths were noticed when walking on sand ($0.026\pm 0.021$) compared with stable ground ($0.023\pm 0.017$) ($p<0.05$; $d=0.15$). This gait adaptation is likely due to the deformable nature of the sandy terrain. Walking on sand showed a significant increase in both swing time and stance time compared to walking on solid ground. The swing time for walking on sand ($0.46\pm 0.057$ s) was significantly shorter than that on solid ground ($0.48\pm 0.056$ s) ($p<0.05$; $d=0.26$). Similarly, the stance time for walking on sand ($0.74\pm 0.09$ s) was significantly longer than that on solid ground ($0.67\pm 0.09$ s) ($p<0.05$; $d=0.57$). Despite these increases in swing and stance times, participants tended to walk faster on sand than on solid ground. The ratio between normalized average walking speed on solid ground ($0.60\pm 0.11$) versus that on sand ($0.64\pm 0.08$) was about $0.94$ ($p<0.05$; $d=0.54$). The mean COM vertical variation during each step on the sand was also observed to increase by $6$%, although this difference was not statistically significant ($p>0.05$; $d=0.01$). This is possibly due to the extra effort from the joints when participants walked on the sand. The gait phase percentages reveal slight differences between walking on sand and solid ground, with the stance phase taking up $42\%$ of the gait cycle on solid ground and $38\%$ on sand.
Differences in joint angles, particularly in the lower extremities, were observed, indicating a distinct adjustment in walking strategy on sand compared to solid ground. Figure 4 shows the profile comparison for the hip, knee, and ankle joint angles throughout the gait cycle. Although there were no clear differences in the overall joint angle patterns, statistical analysis revealed significant differences in peak joint angles. The peak hip flexion angle was significantly greater on sand ($29.36\pm 3.84^{\circ}$) than on solid ground ($p<0.05$; $d=0.11$). Similarly, the peak knee flexion angle was significantly greater on sand ($75.21\pm 7.96^{\circ}$) compared to solid ground ($p<0.05$; $d=0.37$). The peak ankle dorsiflexion angle was also significantly greater on sand ($18.29\pm 7.50^{\circ}$) than on solid ground ($p<0.05$; $d=0.16$). These findings highlight the specific adaptations in lower limb joint kinematics when walking on sand compared to solid ground.
@@ 3.2 Kinetic Analysis Results
Figures 5 and 5 show the longitudinal and vertical GRFs normalized by the participant’s weight, respectively. The shaded region indicated the one-standard variation. It is observed that the normalized magnitude of the maximum longitudinal force $F_{x}$ on sand is relatively smaller than that on solid ground in both forward ($0.0464\pm 0.03$ vs. $0.15\pm 0.035$) and backward directions ($-0.076\pm 0.024$ vs. $-0.1\pm 0.047$). However, regarding the vertical force $F_{z}$, sandy terrain might provide more supporting force during the heel strike ($1.22\pm 0.22$ vs. $1.03\pm 0.084$) compared with the solid ground. It is also noticeable that there is a slight difference with the double-hump pattern of the vertical force $F_{z}$ throughout the contact phase. The magnitudes of two humps on the solid ground appear almost equivalent. Nevertheless, for the sandy terrain, the first hump of the force is slightly higher than the second one in the late contact phase. Similar results were also reported on the level sand in [10]. The significant differences in longitudinal ($p<0.05$; $d=0.58$) and vertical ($p<0.05$; $d=0.36$) GRFs between sand and solid ground indicate that the type of terrain has a substantial impact on the forces exerted by the foot during walking. The increased variability and different magnitudes of GRFs on sand compared to solid ground suggest that walking on sand requires different biomechanical strategies, possibly due to the less stiff surface.
@@ 4.1 Kinematics Features
Different from the experiment setup designed in [3], the participants in this study were asked to walk across two different terrains sequentially from the solid ground to the sand box, allowing each participant to keep the preferred comfortable walking pace. Our findings suggest that walking strategies adapt significantly when transitioning from solid ground to sand. The extended stride length and width observed on the sand were likely compensatory mechanisms to maintain balance on the yielding surface. Participants demonstrated a higher average walking speed on sand compared to solid ground; see the velocity ratio result in Fig. 3. This result is consistent with [7], which reported that the self-selected speed on a $60$ m distance increased when walking on sand with respect to solid ground, with the highest speed on wet sand. This observation suggests that individuals might compensate for the challenging terrain by adopting a faster pace, potentially to minimize the time spent on an unstable surface. Additionally, the increased COM vertical variation indicates an adaptive response to optimize energy efficiency and maintain stability on the yielding terrain, which was also found for the mid-stance phase ($21$-$60$% gait cycle) in [3]. These adaptations are pivotal for navigating through granular substrates, where it is more challenging than navigating on solid ground.
@@
The joint angles, especially in the lower extremities, demonstrate distinct gait variations on solid and sand terrains. Walking on the sandy terrain results in greater dorsiflexion of the ankle in the early stance phase but a similar peak amplitude of the ankle plantar flexion near toe-off. This is because the heel would experience an intrusion into the sand and then the forefoot was prevented from plantarflexion at the beginning of the stance phase; see the small portion of ankle plantarflexion on the solid ground shown as the blue dashed line in Fig. 4. Furthermore, a relatively larger variability of the ankle dorsi/plantarflexion on the sand was found compared with the hip flexion/extension and knee flex/extension. This represents the nature of complexity and instability of deformable terrain locomotion.
The greater flexion observed in the hip and knee joints during the toe-off and swing phases on sand can be interpreted as a compensatory mechanism for the loss of momentum during the stance phase. The deformable nature of sand reduces the efficiency of energy transfer, necessitating additional joint flexion to facilitate foot clearance and forward progression. These kinematic findings underscore the adaptability and fine-tuned control required for efficient and stable locomotion on deformable surfaces. Understanding these adaptations can inform the design of assistive devices, prosthetics, and rehabilitation protocols, as well as guide the development of robotic systems and control algorithms that can effectively traverse challenging terrains.
4.2 Kinetics Characteristics